# Learning Massive-scale Partial Correlation Networks in Clinical Multi-omics Studies with HP-ACCORD

### v2

@@ 3.1 ACCORD loss and estimator
$L(\boldsymbol{\Omega},x)=-\log\det\boldsymbol{\Omega}_{D}+(1/2)\operatorname{\bf tr}(x^{T}\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}x).$ (3.1)
The associated risk $R(\boldsymbol{\Omega})=\operatorname{\bf E}L(\boldsymbol{\Omega},X)=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{\Sigma}^{*})$ is minimized by $\boldsymbol{\Omega}^{*}={\boldsymbol{\Theta}^{*}_{D}}^{-1/2}\boldsymbol{\Theta}^{*}$ since it is the unique solution to the equation $\nabla R(\boldsymbol{\Omega})=\mathbf{0}$ and $R$ is convex. The essential domain of $R$ on which $R$ is finite is $\operatorname{\bf dom}R=\{(m_{ij})\in\mathbb{R}^{p\times p}:m_{ii}>0,~{}i=1,\dotsc,p\}\cap\mathbb{R}^{p\times p}$. Then transformation $T:\boldsymbol{\Theta}\mapsto\boldsymbol{\Theta}_{D}^{-1/2}\boldsymbol{\Theta}$ from $\operatorname{\bf dom}R$ to itself is bijective with inverse $T^{-1}:\boldsymbol{\Omega}\mapsto\boldsymbol{\Omega}_{D}\boldsymbol{\Omega}$. Obviously $\boldsymbol{\Omega}^{*}=T(\boldsymbol{\Theta}^{*})$. In other words, the population risk is uniquely minimized by a one-to-one transformation $T$ of the true precision matrix $\boldsymbol{\Theta}^{*}$. If $\hat{\boldsymbol{\Omega}}$ minimizes the sample average of (3.1), then it consistently estimates $\boldsymbol{\Omega}^{*}$. Hence $T^{-1}(\hat{\boldsymbol{\Omega}})$ consistently estimates $\boldsymbol{\Theta}^{*}=T^{-1}(\boldsymbol{\Omega}^{*})$, as long as the dimension $p$ is fixed.
@@
$\hat{\boldsymbol{\Omega}}=\operatorname*{arg\,min}\nolimits_{\boldsymbol{\Omega}\in\mathbb{R}^{p\times p}}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\lambda\|\boldsymbol{\Omega}\|_{1}\right\};$ (3.2)
recall that $\boldsymbol{S}$ is the sample covariance matrix. An important point here is that the transformation pair $(T,T^{-1})$ preserves the sparsity pattern of their respective arguments: if $\omega_{ij}=[T(\boldsymbol{\Theta})]_{ij}$ with $\boldsymbol{\Theta}=(\theta_{ij})$, then $\omega_{ij}=\theta_{ij}/\sqrt{\theta_{ii}}$ and $\theta_{ij}=\omega_{ii}\omega_{ij}$. Therefore, $\omega_{ij}=0$ if and only if $\theta_{ij}=0$, and $\boldsymbol{\Omega}^{*}$ shares the sparsity pattern with $\boldsymbol{\Theta}^{*}$. The $\ell_{1}$ penalization on $\boldsymbol{\Omega}$ in (3.2) is thus justified. Since the downstream task is to analyze the partial correlations ($\rho^{*}_{ij})$, using the relation $\rho^{*}_{ij}=-\theta^{*}_{ij}/\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$ and $\theta^{*}_{ij}=\omega^{*}_{ii}\omega^{*}_{ij}$, we can consistently estimate $\rho^{*}_{ij}$ by the formula $\hat{\rho}_{ij}=-(1/2)(\hat{\omega}_{ij}/\hat{\omega}_{jj}+\hat{\omega}_{ji}/\hat{\omega}_{ii})$, in which $\hat{\boldsymbol{\Omega}}=(\hat{\omega}_{ij})$.
The loss (3.1) and estimator (3.2) resemble those of CONCORD in §2. The critical difference is that the argument $\boldsymbol{\Omega}$ is allowed to be asymmetric in our proposal. This simple change makes the estimator consistent (at least in the low-dimensional regime; high-dimensional results are presented in §4), a feature absent in CONCORD. For an obvious reason, we name the loss function (3.1) the ACCORD loss, and estimator (3.2) the ACCORD estimator, after Asymmetric ConCORD.
@@ 3.2 ACCORD-FBS algorithm
Assume the condition for Theorem 3.1, and that the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ is generated by Algorithm 1 with the step size sequence $\{\tau_{t}\}$ satisfying either i) $\beta=1$, $\tau_{0}\in[\underline{\tau},\bar{\tau}]$, where $0<\underline{\tau}\leq\bar{\tau}<2/L$, or ii) $10<\beta<$ 1. Then, the objective value sequence $\{f(\boldsymbol{\Omega}^{(t)})\}$ converges to the minimum $f^{\star}=f(\hat{\boldsymbol{\Omega}})$ monotonically. Furthermore, if the initial iterate $\boldsymbol{\Omega}^{(0)}$ is chosen such that $f(\boldsymbol{\Omega}^{(0)})\leq\operatorname{\bf tr}(\boldsymbol{S})/2+\lambda p$, then the following holds
@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
For its simplicity and linear convergence rate, the ACCORD-FBS algorithm (Algorithm 1) has an advantage in scaling up to handle massive-scale data. Note that the main computational components of ACCORD-FBS are: 1) sparse-dense matrix multiplication in computing the gradient $\nabla g(\boldsymbol{\Omega}^{(t)})=\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$ and 2) element-wise operations in computing (3.5); note that $\boldsymbol{\Omega}^{(t)}$ is sparse by construction. The latter is “embarassingly parallel.” The former can also be easily parallelized in shared memory systems such as those employing graphical processing units (GPUs). However, if the size of the data becomes massive so that it does not fit into the system memory, employment of a distributed memory system becomes necessary. In distributed computation, communication cost becomes a major factor of the performance.
Our distributed-memory HPC implementation of ACCORD-FBS, termed HP-ACCORD, iteratively conducts the sparse-dense matrix multiplication for the gradient step as a special case of the SpDM3 algorithm [20]. This algorithm, named one-dimensionally distributed matrix multiplication (1DMM), is summarized in Algorithm 2. In HP-ACCORD, $\boldsymbol{\Omega}$ and $\boldsymbol{S}$ are separated by columns (row-wise separation is also possible for $\boldsymbol{\Omega}$) and store each partition in different computational nodes. Then, the gradient $\nabla g(\boldsymbol{\Omega})$ is computed with 1DMM. While the standard method for distributed matrix-matrix multiplication on HPC systems is the scalable universal matrix multiplication algorithm [SUMMA, 36], this method partitions involved matrices by both rows and columns and allocates them across computational nodes, and completing each block of the product requires multiple rounds of broadcasting submatrices of both operands, which results in a major performance bottleneck. On the other hand, 1DMM sends and receives one-dimensional blocks of only one operand, i.e., $\boldsymbol{\Omega}$ in Algorithm 2. Because of this difference, the communication cost of Algorithm 2 is smaller than SUMMA since $\boldsymbol{\Omega}$ is sparse while $\boldsymbol{S}$ is dense [19]. Computation of the gradient can alternatively be conducted in two steps: $\boldsymbol{Y}=\boldsymbol{\Omega}\boldsymbol{X}^{T}$ and $\nabla g(\boldsymbol{\Omega})=(1/n)\boldsymbol{Y}\boldsymbol{X}$, each of which can be computed with 1DMM. This strategy is advantageous when $n$ is much smaller than $p$ [20].
@@ 3.4 Tuning
The choice of the regularization parameter $\lambda$ impacts the practical performance of the ACCORD estimator. At the omics scale, sample reuse methods such as cross validation or neighborhood selection [28] are ruled out since they incur a nontrivial number of expensive passes of computing the estimator. Following CONCORD and SPACE, which are also pseudolikelihood-based methods, we adopt a Bayesian information criterion (BIC)-type approach for tuning. Specifically, using the loss function part of (3.2), we choose $\lambda$ minimizing
@@ 3.5 Bias correction
$\tilde{\boldsymbol{\Omega}}=\operatorname*{arg\,min}_{\boldsymbol{\Omega}:\;\boldsymbol{\Omega}_{\hat{S}_{\lambda}^{c}}=0}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\phi\lambda\|\boldsymbol{\Omega}\|_{1}\right\},$ (3.8)
where $0\leq\phi\leq 1$. The refitted estimator $\tilde{\boldsymbol{\Omega}}=(\tilde{\omega}_{ij})$ can be computed efficiently using Algorithm 1, by replacing the $\lambda$ with $\phi\lambda$ for $(i,j)\in\hat{S}_{\lambda}$ and with $\infty$ for $(i,j)\not\in\hat{S}_{\lambda}$.
@@ 4 Statistical Properties
In this section, we show that the ACCORD estimator can consistently estimate $\boldsymbol{\Omega}^{\ast}$, a one-to-one reparameterization of the true precision matrix $\boldsymbol{\Theta}^{*}$ in various measures, under appropriate conditions. All the results provided here are non-asymptotic. In addition to the vector $\ell_{1}$ and $\ell_{\infty}$ norms introduced in §1, we use $\vvvert\boldsymbol{M}\vvvert$ to denote an operator norm of matrix $\boldsymbol{M}=(m_{ij})$ induced by the underlying vector norm. In particular, $\vvvert\boldsymbol{M}\vvvert_{\infty}=\max_{\boldsymbol{x}\neq 0}\frac{\|\boldsymbol{M}\boldsymbol{x}\|_{\infty}}{\|\boldsymbol{x}\|_{\infty}}=\max_{i}\sum_{j=1}^{p}|m_{ij}|$. The Frobenius norm $\|\boldsymbol{M}\|_{F}$ of $\boldsymbol{M}$ is its vector $\ell_{2}$ norm. For an $l\times m$ matrix $\boldsymbol{M}$, $A\subset[l]:=\{1,\dotsc,l\}$, and $B\subset[m]$, we denote by $\boldsymbol{M}_{AB}$ the $|A|\times|B|$ submatrix of $\boldsymbol{M}$ taking the rows and columns of $\boldsymbol{M}$ with indices in $A$ and $B$, respectively.
@@ 4.1 Estimation error bounds
The consistency of $\hat{\boldsymbol{\Omega}}$ in terms of $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$ is obtained if $\sigma^{2}$ and $\kappa_{\Omega^{\ast}}$ are bounded, at the rate of $O_{P}(\sqrt{n^{-1}|S|\log p})$. The latter is bounded if, e.g., the graph implied by $\boldsymbol{\Theta}^{*}$ has a finite maximum degree of $d=\max_{i\in[p]}|\{j\in[p]:\allowbreak\theta^{\ast}_{ij}\neq 0\}|$: since $|\theta^{*}_{ij}|\leq\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$, $\kappa_{\Omega^{\ast}}=\max_{i\in[p]}\sum_{j=1}^{p}|\theta^{*}_{ij}|/\sqrt{\theta^{*}_{ii}}\leq d\max_{i\in[p]}\sqrt{\theta^{*}_{ii}}\leq d/\sqrt{\lambda_{\min}(\boldsymbol{\Sigma}^{*})}$, where $\lambda_{\min}(\boldsymbol{\Sigma}^{*})\geq\alpha$ is the minimum eigenvalue of $\boldsymbol{\Sigma}^{*}$. This rate and the sample complexity of $n\gtrsim|S|\log p$ match those for the graphical lasso [39, Proposition 11.9].

### v3

@@ 3.1 ACCORD loss and estimator
$L(\boldsymbol{\Omega},x)=-\log\det\boldsymbol{\Omega}_{D}+(1/2)\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}xx^{T}).$ (3.1)
For the associated risk $R(\boldsymbol{\Omega}):=\operatorname{\bf E}L(\boldsymbol{\Omega},X)=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{\Sigma}^{*})$ with the essential domain $\operatorname{\bf dom}R=\{(m_{ij})\in\mathbb{R}^{p\times p}:m_{ii}>0,~{}i=1,\dotsc,p\}\cap\mathbb{R}^{p\times p}$ on which $R$ is finite, the following theorem holds (details are in the Supplemental Material).
Theorem 3.1.
The ACCORD risk $R(\boldsymbol{\Omega})$ is uniquely minimized by $\boldsymbol{\Omega}^{*}:={\boldsymbol{\Theta}^{*}_{D}}^{-1/2}\boldsymbol{\Theta}^{*}$.
In light of Theorem 3.1, define the transformation
$T:\boldsymbol{\Theta}\mapsto\boldsymbol{\Theta}_{D}^{-1/2}\boldsymbol{\Theta}$
from $\operatorname{\bf dom}R$ to itself. One important property of $T$ is that it is continuous and bijective with inverse $T^{-1}:\boldsymbol{\Omega}\mapsto\boldsymbol{\Omega}_{D}\boldsymbol{\Omega}$. In other words, the population risk is uniquely minimized by a one-to-one transformation $T$ of the true precision matrix $\boldsymbol{\Theta}^{*}$. In the sample version of $R$ where the sample covariance matrix $\boldsymbol{S}$ replaces the $\boldsymbol{\Sigma}^{*}$, the unique minimizer of the empirical risk $\hat{R}_{n}(\boldsymbol{\Omega})=-\log\det\boldsymbol{\Omega}_{D}+\frac{1}{2}\operatorname{\bf tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})$ is given by
$\hat{\boldsymbol{\Omega}}=T(\tilde{\boldsymbol{\Theta}})=\tilde{\boldsymbol{\Theta}}_{D}^{-1/2}\tilde{\boldsymbol{\Theta}},\quad\tilde{\boldsymbol{\Theta}}=\boldsymbol{S}^{-1}.$
in the low-dimensional regime where $\boldsymbol{S}$ is positive definite. Then, $\hat{\boldsymbol{\Omega}}\to\boldsymbol{\Omega}^{*}$ as $\boldsymbol{S}\to\boldsymbol{\Sigma}^{*}$ by the continuous mapping theorem. Also, if we let $\hat{\boldsymbol{\Theta}}=T^{-1}(\hat{\boldsymbol{\Omega}})=\hat{\boldsymbol{\Omega}}_{D}\hat{\boldsymbol{\Omega}}$, then $\hat{\boldsymbol{\Theta}}\to\boldsymbol{\Theta}^{*}$. Likewise, the partial correlations $(\rho^{*}_{ij})$ can be consistently estimated with $\hat{\rho}_{ij}=-(1/2)(\hat{\omega}_{ij}/\hat{\omega}_{jj}+\hat{\omega}_{ji}/\hat{\omega}_{ii})$ when $\hat{\boldsymbol{\Omega}}=(\hat{\omega}_{ij})$, using the relation $\rho^{*}_{ij}=-\theta^{*}_{ij}/\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$ and $\theta^{*}_{ij}=\omega^{*}_{ii}\omega^{*}_{ij}$.
@@
$\hat{\boldsymbol{\Omega}}=\operatorname*{arg\,min}\nolimits_{\boldsymbol{\Omega}\in\mathbb{R}^{p\times p}}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\lambda\|\boldsymbol{\Omega}\|_{1}\right\}.$ (3.2)
The $\ell_{1}$ penalization on $\boldsymbol{\Omega}$ in (3.2) is justified by another important property that $T$ and $T^{-1}$ preserve the support of their arguments; if $\omega_{ij}=[T(\boldsymbol{\Theta})]_{ij}$ with $\boldsymbol{\Theta}=(\theta_{ij})$, then $\omega_{ij}=\theta_{ij}/\sqrt{\theta_{ii}}$ and $\theta_{ij}=\omega_{ii}\omega_{ij}$. Therefore, $\omega_{ij}=0$ if and only if $\theta_{ij}=0$, and hence, $\boldsymbol{\Omega}^{*}$ and $\boldsymbol{\Theta}^{*}$ share the same sparsity pattern.
The loss (3.1) and estimator (3.2) resemble those of CONCORD in §2. The critical difference is that the argument $\boldsymbol{\Omega}$ is allowed to be asymmetric in our proposal. This simple change makes the estimator consistent and recover the support of $\boldsymbol{\Theta}^{*}$ (at least in the low-dimensional regime; high-dimensional results are presented in §4), a feature absent in CONCORD. (The latter estimates a root of (2.4). Let alone uniqueness, the existence of its closed form expression is unclear, and it is unlikely that it preserves the sparsity pattern of $\boldsymbol{\Theta}^{*}$.) Due to the asymmetric nature of our approach, we name the loss function (3.1) the ACCORD loss, and the estimator (3.2) the ACCORD estimator, after Asymmetric ConCORD.
@@ 3.2 ACCORD-FBS algorithm
Assume the condition for Theorem 3.2, and that the iterate sequence $\{\boldsymbol{\Omega}^{(t)}\}$ is generated by Algorithm 1 with the step size sequence $\{\tau_{t}\}$ satisfying either i) $\beta=1$, $\tau_{0}\in[\underline{\tau},\bar{\tau}]$, where $0<\underline{\tau}\leq\bar{\tau}<2/L$, or ii) $10<\beta<$ 1. Then, the objective value sequence $\{f(\boldsymbol{\Omega}^{(t)})\}$ converges to the minimum $f^{\star}=f(\hat{\boldsymbol{\Omega}})$ monotonically. Furthermore, if the initial iterate $\boldsymbol{\Omega}^{(0)}$ is chosen such that $f(\boldsymbol{\Omega}^{(0)})\leq f(\boldsymbol{I}_{p})=\operatorname{\bf tr}(\boldsymbol{S})/2+\lambda p$, then the following holds.
@@
Remark 3.1.
In Algorithm 1, the computational complexity of each iteration is $O(np^{2})$, with the bottleneck being the gradient computation $\nabla g(\boldsymbol{\Omega}^{(t)})\leftarrow(1/n)(\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T})\boldsymbol{X}$; see Appendix C for comparison of the complexity with other methods. In fact, the required number of arithmetic operations can be further reduced by exploiting the sparsity of the iterate $\boldsymbol{\Omega}^{(t)}$. Moreover, matrix multiplication of the type $\boldsymbol{\Omega}^{(t)}\boldsymbol{X}^{T}$ can be distributed among multiple computational nodes, allowing the algorithm to be more efficient in HPC environments.
@@ 3.3 HP-ACCORD: HPC implementation of ACCORD
For its simplicity and linear convergence rate, the ACCORD-FBS algorithm (Algorithm 1) has an advantage in scaling up to handle massive-scale data. Note that the main computational components of ACCORD-FBS are: 1) sparse-dense matrix multiplication in computing the gradient $\nabla g(\boldsymbol{\Omega}^{(t)})=\boldsymbol{\Omega}^{(t)}\boldsymbol{S}$ and 2) element-wise operations in computing (3.5); note that $\boldsymbol{\Omega}^{(t)}$ is sparse by construction. The latter is “embarassingly parallel.” The former can also be easily parallelized in shared-memory systems, such as those employing graphical processing units (GPUs). However, if the size of the data becomes massive so that it does not fit into the system memory, employment of a distributed memory system becomes necessary. In distributed computation, communication cost becomes a significant factor of the performance.
Our distributed-memory HPC implementation of ACCORD-FBS, termed HP-ACCORD, iteratively conducts the sparse-dense matrix multiplication for the gradient step as a special case of the SpDM3 algorithm [21]. This algorithm, named one-dimensionally distributed matrix multiplication (1DMM), is summarized in Algorithm 2. In HP-ACCORD, $\boldsymbol{\Omega}$ and $\boldsymbol{S}$ are separated by columns (row-wise separation is also possible for $\boldsymbol{\Omega}$) and store each partition in different computational nodes. Then, the gradient $\nabla g(\boldsymbol{\Omega})$ is computed with 1DMM. While the standard method for distributed matrix-matrix multiplication on HPC systems is the scalable universal matrix multiplication algorithm [SUMMA, 39], this method partitions involved matrices by both rows and columns and allocates them across computational nodes, and completing each block of the product requires multiple rounds of broadcasting submatrices of both operands, which results in a significant performance bottleneck. On the other hand, 1DMM sends and receives one-dimensional blocks of only one operand, i.e., $\boldsymbol{\Omega}$ in Algorithm 2. Because of this difference, the communication cost of Algorithm 2 is smaller than SUMMA since $\boldsymbol{\Omega}$ is sparse while $\boldsymbol{S}$ is dense [20]. Computation of the gradient can alternatively be conducted in two steps: $\boldsymbol{Y}=\boldsymbol{\Omega}\boldsymbol{X}^{T}$ and $\nabla g(\boldsymbol{\Omega})=(1/n)\boldsymbol{Y}\boldsymbol{X}$, each of which can be computed with 1DMM. This strategy is advantageous when $n$ is much smaller than $p$ [21].
@@ 3.4 Tuning
The choice of the regularization parameter $\lambda$ impacts the practical performance of the ACCORD estimator. At the omics scale, sample reuse methods such as cross-validation or neighborhood selection [28] are ruled out, as they incur a nontrivial number of expensive passes to compute the estimator. Following CONCORD and SPACE, which are also pseudolikelihood-based methods, we adopt a Bayesian information criterion (BIC)-type approach for tuning. Specifically, using the loss function part of (3.2), we choose $\lambda$ minimizing
@@ 3.5 Bias correction
$\breve{\boldsymbol{\Omega}}=\operatorname*{arg\,min}_{\boldsymbol{\Omega}:\;\boldsymbol{\Omega}_{\hat{S}_{\lambda}^{c}}=0}\left\{-\log\det\boldsymbol{\Omega}_{D}+(1/2)\text{tr}(\boldsymbol{\Omega}^{T}\boldsymbol{\Omega}\boldsymbol{S})+\phi\lambda\|\boldsymbol{\Omega}\|_{1}\right\},$ (3.8)
where $0\leq\phi\leq 1$. The refitted estimator $\breve{\boldsymbol{\Omega}}=(\breve{\omega}_{ij})$ can be computed efficiently using Algorithm 1, by replacing the $\lambda$ with $\phi\lambda$ for $(i,j)\in\hat{S}_{\lambda}$ and with $\infty$ for $(i,j)\not\in\hat{S}_{\lambda}$.
@@ 4 Statistical Properties
In this section, we show that the ACCORD estimator can consistently estimate $\boldsymbol{\Omega}^{\ast}$, a one-to-one reparameterization of the true precision matrix $\boldsymbol{\Theta}^{*}$ in various measures, under appropriate conditions. All the results provided here are non-asymptotic. In addition to the vector $\ell_{1}$ and $\ell_{\infty}$ norms introduced in §1, we use $\vvvert\boldsymbol{M}\vvvert$ to denote an operator norm of matrix $\boldsymbol{M}=(m_{ij})$ induced by the underlying vector norm. In particular, $\vvvert\boldsymbol{M}\vvvert_{\infty}=\max_{\boldsymbol{x}\neq 0}\frac{\|\boldsymbol{M}\boldsymbol{x}\|_{\infty}}{\|\boldsymbol{x}\|_{\infty}}=\max_{i}\sum_{j=1}^{p}|m_{ij}|$. The Frobenius norm $\|\boldsymbol{M}\|_{F}$ of $\boldsymbol{M}$ is its vector $\ell_{2}$ norm. For a finite set $A$, we denote by $|A|$ the number of elements in $A$. For an $l\times m$ matrix $\boldsymbol{M}$, $A\subset[l]:=\{1,\dotsc,l\}$, and $B\subset[m]$, we denote by $\boldsymbol{M}_{AB}$ the $|A|\times|B|$ submatrix of $\boldsymbol{M}$ taking the rows and columns of $\boldsymbol{M}$ with indices in $A$ and $B$, respectively.
@@ 4.1 Estimation error bounds
The consistency of $\hat{\boldsymbol{\Omega}}$ in terms of $\|\cdot\|_{1}$ and $\|\cdot\|_{F}$ is obtained if $\sigma^{2}$ and $\kappa_{\Omega^{\ast}}$ are bounded, at the rate of $O_{P}(\sqrt{n^{-1}|S|\log p})$. The latter can be bounded, e.g., when the graph implied by $\boldsymbol{\Theta}^{*}$ has a bounded number of non-zero entries per row $d=\max_{i\in[p]}|\{j\in[p]:\allowbreak\theta^{\ast}_{ij}\neq 0\}|$. Then $|\theta^{*}_{ij}|\leq\sqrt{\theta^{*}_{ii}\theta^{*}_{jj}}$ and $\kappa_{\Omega^{\ast}}=\max_{i\in[p]}\sum_{j=1}^{p}|\theta^{*}_{ij}|/\sqrt{\theta^{*}_{ii}}\leq d\max_{i\in[p]}\sqrt{\theta^{*}_{ii}}\leq d/\sqrt{\lambda_{\min}(\boldsymbol{\Sigma}^{*})}$, where $\lambda_{\min}(\boldsymbol{\Sigma}^{*})\geq\alpha$ is the minimum eigenvalue of $\boldsymbol{\Sigma}^{*}$.This rate and the sample complexity of $n\gtrsim|S|\log p$ match those for the graphical lasso [42, Proposition 11.9].